If P(k) = k2

(k + 2)(k – 1) is true, then what is P (k + 1)? 


1
Expert's answer
2021-05-07T14:22:13-0400

Solution:

"P(k)=k^2(k+2)(k-1)"

Now, replace "k" with "(k+1)",

"P(k+1)=(k+1)^2(k+1+2)(k+1-1)
\\=(k^2+1+2k)(k+3)(k)
\\=k(k+3)(k^2+2k+1)"